OpenStudy (anonymous):

If f(x) = x3, which of the following describes the graph of f(x - 3)? A.) The graph of f(x - 3) is a vertical shift of f(x) = x3 three units up. B.) The graph of f(x - 3) is a vertical shift of f(x) = x3 three units down. C.) The graph of f(x - 3) is a horizontal shift of f(x) = x3 three units to the right. D.) The graph of f(x - 3) is a horizontal shift of f(x) = x3 three units to the left.

OpenStudy (anonymous):

you mean (3)(x) or x^3 ??

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I think is in between B or C i am not so sure

OpenStudy (anonymous):

well yo didn't tell me you mean 3x or x^3 !

OpenStudy (anonymous):

oh sorry is x^3

OpenStudy (anonymous):

so f(x-3) = (x-3)^2 = x^2 + 9 - 6x = f(x) + 9 - 6x

OpenStudy (anonymous):

So its C? or B?

OpenStudy (anonymous):

it's C